Face Value

$500,000

Coupon Rate

5.0%

Market Rate

4.0%

Semiannual Interest Payments

Due June 30 and Dec 31

Maturity Date

5 years

Issue Date

Jan.1

Based on the data above, complete the journal entries for:

  1. The issue of the bonds on January 1
  2. The payment of interest and amortization of the premium on June 30 and
  3. the payment of interest and amortization of the premium on December 31

Date

General Journal

Debit

Credit

1/1

6/30

12/31

Face Value

$200,000

Coupon Rate

6.5%

Market Rate

7.5%

Semiannual Interest Payments

Due June 30 and Dec 31

Maturity Date

5 years

Issue Date

Jan.1

Based on the data above, complete the journal entries for:

  1. The issue of the bonds on January 1
  2. The payment of interest and amortization of the discount on June 30
  3. And the payment of interest and amortization of the discount on December 31

Date

General Journal

Debit

Credit

1/1

6/30

12/31
Answer #1

1)semiannual interest =Par value *coupon rate *n/12    [2semiannual periods in a year comprising of 6 months each]

                             = 500000*.05*6/12

                             = 12500

semiannual market interest rate = 4*6/12 = 2%

semiannual periods = 5*2 = 10

Present value of Bond =[PVA2%,10*semiannual interest ]+[PVF2%,10*par value]

            =8.98259*12500]   +[.82035*500000]

             = 112282.38+ 410175

             = 522457.38    [rounded to 522457]

Date Account title Debit credit
1Jan cash 522457
Premium on bond payable 22457
Bond payable 500000
30June Interest expense (522457*2%) 10449.14
Premium on bond payable 2050.86
cash 12500
31Dec Interest expense (520406.14*2%) 10408.12
Premium on bond payable 2091.88
cash 12500

carrying value of Bond on 30June =Issue price +premium amortized on 30June

                               = 522457-2050.86

                                = 520406.14

2)

semiannual interest =Par value *coupon rate *n/12    [2semiannual periods in a year comprising of 6 months each]

                             = 200000*.065*6/12

                             = 6500

semiannual market interest rate = 7.5*6/12 = 3.75%

semiannual periods = 5*2 = 10

Present value of Bond =[PVA3.75%,10*semiannual interest ]+[PVF3.75%,10*par value]

            =[8.21279*6500]    +[.69202*200000]

             =53383.14+ 138404

             = 191787.14    [rounded to 191787]

Date Account title Debit credit
1Jan cash 191787
Discount on bond payable 8213
Bond payable 200000
30June Interest expense (191787*.0375) 7192.01
Discount on bond payable 692.01
cash 6500
31Dec Interest expense (192479.01*.0375) 7217.96
Discount on bond payable 717.96
cash 6500

carrying value of Bond on 30June =Issue price +premium amortized on 30June

                            = 191787 + 692.01

                       = 192479.01

**Find present value annuity factor and present value table from their table respectively at i% for n periods or using financial calculator .